Check my work 3 Miller Companys contribution format income statement for the most recent month is shown below Total 108,000 108,000 Per Unit s 216,000 6.00 3.00 Sales (36,000 units) Variable expenses Contribution margin Fixed expenses Net operating income 3.00 nts Skipped $ 58,000 Required: Consider each case independently): eBook Hint 1. What is the revised net operating income if unit sales increase by 11%? 2. What is the revised net operating income if the selling price decreases by $1.20 per unit and the number of units sold increases by 23%? 3. What is the revised net operating income if the selling price increases by $1.20 per unit, fixed expenses increase by $8,000, and the number of units sold decreases by 4%? 4, what is the revised net operating income if the selling price per unit increases by 20%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 11%? Print eferences 1. Net operating income 2. Net operating income 3. Net operating income 4. Net operating income
Answer #1

Calculate net operating income :

a) Net operating income = (36000*111%*3)-50000 = $69880

b) Net operating income = (36000*123%*1.8)-50000 = $29704

c) Net operating income = (36000*96%*4.2)-58000 = $87152

d) Net operating income = (36000*89%*3.9)-50000 = $74956